A 29.55-g sample of a solid is placed in a flask. Toluene, in which the solid is insoluble, is added to the flask so that the total volume of solid and liquid together is 50.0 mL. The solid and toluene together weigh 49.71 g. The density of toluene at the temperature of the experiment is 0.864 g/mL. What is the density of the solid?
1
Expert's answer
2018-01-23T15:10:08-0500
We find the mass of toluene, which is added to the flask:
m_toluene=m_total-m_solid=49.71-29.55=20.16 g
Based on the density value for toluene, we find the volume of toluene, which is added to the flask:
V=m/ρ=20.16/0.864=23.33 mL
We find the volume of solid, which is added to the flask:
V_solid=m_total-m_toluene=50-23.33=26.67 mL
Knowing the volume and mass of a solid, we find its density:
ρ=m/V=29.55/26.67=1.11 g/mL
Answer: ρ = 1.11 g/mL.